OpenStudy (jennie):

P (x) = 90(1 + 1.5x) / 1 + 0.5x P = number of fish x = number of months since the fish were placed in the aquarium Six months after the fish are first placed in the aquarium, what is the total number of fish?

jimthompson5910 (jim_thompson5910):

I'm guessing it should look like this \[\Large P(x) = \frac{90(1+1.5x)}{1+0.5x}\]

OpenStudy (jennie):

yes I need to find out how many fish after 1 month, then 3 months, then 6 months and then 12 month

jimthompson5910 (jim_thompson5910):

You do the same for x = 3 months, x = 6 months and x = 12 months (12 months = 1 yr)

jimthompson5910 (jim_thompson5910):

oops made a typo, should be 150 exactly, let me fix it

jimthompson5910 (jim_thompson5910):

\[\Large P(x) = \frac{90(1+1.5x)}{1+0.5x}\] \[\Large P(1) = \frac{90(1+1.5*1)}{1+0.5*1}\] \[\Large P(1) = \frac{90(1+1.5)}{1+0.5}\] \[\Large P(1) = \frac{90(2.5)}{1.5}\] \[\Large P(1) = \frac{225}{1.5}\] \[\Large P(1) = 150\] So after one month, there are 150 fish in the aquarium.